OpenStudy (anonymous):

How do I do this s=ut+1/2at^2 find a when u =15 t=3?

OpenStudy (anonymous):

substitution? do you have any answer choices

OpenStudy (anonymous):

you need to substitution first..

OpenStudy (anonymous):

@Nurali

OpenStudy (anonymous):

s = 15(3) +0.5 a (9) a = 2 (s-45)/9 Need s to solve this.